EXTCONV(1) EXTCONV(1) NAME extconv−convertaTeXdocumentineitherBig5+orGBKencodinginto‘preprocessed’form. SYNOPSIS extconv<infile>outfile DESCRIPTION Big5+andGBKencodingsusethecharacters{,},and\whichhavespecialmeaningsinTEX documents. AfterprocessingaLATEX2εdocumentwhichcontainssuchcharacterswithextconvyouneednot careaboutthesespecialcharacters. ThisfilterispartoftheCJKmacropackageforLATEX2ε. SEEALSO bg5conv(1),sjisconv(1),cefconv(1),cef5conv(1),cefsconv(1), theCJKdocumentationfiles. AUTHOR WernerLemberg<wl@gnu.org> CJKVersion4.3.0 20-Jun-1999 1